Grade 9Algebra

If [(x²-y²)/(x²+y²)]+[(x²+y²)/(x²-y²)]=√2,then find the value of (x8-y8)/(x8+y8)

Amlan Sahoo
8 Years agoGrade 9

1 Answer

Arun
8 Years ago
Dear student
 
In given part
Taking LCM and then solving
 
[(x²-y²)² + (x²+y²)² ]/ (x⁴-y⁴) = sqrt(2)
 
2(x⁴ + y⁴)/ (x⁴ - y⁴) = sqrt(2)
 
(x⁴ + y⁴) /( x⁴ - y⁴) = 1/ sqrt(2)
 
Now you can solve it
 
 
 
Regards
Arun (askIITians forum expert)